Anita has a debt of RM5,000 to pay in 9 months. Anita received a new agreement to make the two equal payment, one now and another at the end of the 18 months. Find the size of this payment if interest is charged 9% per annum using the 9 months as the focal date.
1
Expert's answer
2021-09-10T08:09:55-0400

Assume repayment =x

left debt =5000-x (after the first repayment)

after 18 months we have "5000-x\times 9\%/12\times18=(5000-x)\times 13.5\% interest"

"X=(5000-x)\times 13\% +(5000-x)"

"x=2658.1"

the repayment is Rm2658.1